OpenStudy (anonymous):

Factor each difference of two squares 4x^2-25 and show work.

OpenStudy (nannie):

4x^2-25 -25 + 25 + 4x2 = 0 + 25 0 + 4x2 = 0 + 25 4x2 = 0 + 25 4x2 = 25 Divide by '4'. x2 = 6.25 x = {-2.5, 2.5}